# Scaling Properties of Speech Language Models

### v1

@@ 1 Introduction
  • We speculate that SLMs require more context than fits in their context window to acquire from commonly used speech datasets the semantic understanding measured by our metrics. Accordingly, we propose a new speech dataset to boost semantic understanding in SLMs. Specifically, we synthesized a spoken version of the Tiny Stories dataset Eldan and Li (2023), and show that its use during pre-training improves semantic downstream performance.
@@ 3.3.2 Data budgets
In order to have a representative set of samples to fit Equations 4 and 7, for each model size, we performed training runs with a ratio of training tokens $D$ to parameters $N$: $D/N\in\{2,4,8,16,32,64,128\}$. This setup yields single-epoch and multi-epoch runs for the larger models (e.g. for the model with 155M parameters the maximum number of training tokens corresponds to 1.82 epochs), but not for the smaller models (e.g. for the model with 85M parameters the maximum number of training tokens corresponds to 0.99 epochs). To have better samples to fit Equation 7, we did additional experiments so that for each model size there were runs with training epochs in $\{2,4,8,10\}$.
@@ 4.2.1 Downstream scaling with compute
We analyze the relationship between the upstream and linguistic downstream performance in SLMs and LLMs. Figure 4 shows the obtained results. Syntactic and semantic downstream metrics before saturation are strongly correlated with the upstream test loss in both LLMs and SLMs. Therefore, the envelope of maximum downstream performance per FLOP also follows a power law, i.e. for a downstream performance function $Q$, $Q\propto C^{\gamma_{q}}$. The power laws for the different performance metrics are depicted in Figure 2 and the exponents presented in Table 3.
@@ 5 Related work
This paper is perhaps most closely related to the work of Hassid et al. (2023). We largely follow their setup in terms of model architecture and evaluation metrics. Similar to our work, they reported improved linguistic downstream performance with scale in SLMs. However, they did not characterize their scaling behavior. Our models of the scaling behavior of SLMs allow practitioners to determine the compute needed to attain a specific loss, syntactic and/or semantic performance; and the optimal allocation of that compute with respect to parameters and tokens. To the best of our knowledge, we are the first to model the scaling properties of downstream linguistic performance in SLMs, and to study the scaling of the considered downstream metrics on text-based LLMs. This enables a comparison between the two modalities in terms of scaling efficiency.
@@ 6.1 On our results
6.2 On limitations
Any extrapolation from our models of the scaling behavior of SLMs should be considered optimistic for the following reasons:
  1. 1.
    The LLMs from the Pythia suite that we used in this study are likely overtrained (all models were trained with $\sim$300B tokens). Optimally trained LLMs (according to Equation 6) should show better performance with scale, and therefore widen the gap with the scaling efficiency of SLMs.
  2. 2.
    The envelope of minimal loss per FLOP (Figure 1) might show a slight negative curvature at larger scale Hoffmann et al. (2022), reducing the scaling efficiency.
  3. 3.
    Our models for downstream performance ignore the fact that the metrics saturate. As observed in text LLMs, the improvements with scale slow down as performance approaches the saturation value. It is likely that, due to saturation, the compute required to yield a particular performance will be larger than predicted. Moreover, due to the lower density of linguistic information per context window in SLMs relative to LLMs, the saturation values of the metrics may be lower for SLMs.
@@ 7 Conclusions
We have trained a large set of SLMs of different sizes and on different data budgets. Using the collected data from those experiments, we studied the scaling properties of their upstream and downstream performance using recently proposed models of scaling laws for neural LMs. The obtained models of the scaling behavior allow practioners to optimally allocate compute to attain a specific loss, syntactic and/or semantic performance. We showed that the pre-training loss and downstream linguistic performance of SLMs and LLMs is highly correlated, and that they both scale predictably according to power laws. This predictable behavior allowed us to compare the scaling properties of SLMs and LLMs, from which we established that the linguistic abilities of SLMs scale up to three orders of magnitude more slowly than those of LLMs. Additionally, we proposed a new speech dataset, sTinyStories, and showed that its use during pre-training improves downstream semantic performance in SLMs. Finally, we explored the use of coarser speech tokenizations as a method to increase the ability of SLMs to model long-range dependencies. However, our results suggest that this is detrimental to downstream performance.

### v2

@@ 1 Introduction
  • We speculate that SLMs require more context than fits in their context window to acquire from commonly used speech datasets the semantic understanding measured by our metrics. Accordingly, we propose a new speech dataset to boost semantic understanding in SLMs. Specifically, we synthesized a spoken version of the Tiny Stories dataset Eldan and Li (2023), and show that its use during pre-training improves downstream semantic performance.
@@ 3.3.2 Data budgets
In order to have a representative set of samples to fit Equations 4 and 7, for each model size, we performed training runs with a ratio of training tokens $D$ to parameters $N$: $D/N\in\{2,4,8,10,20,32,64,100\}$. This setup yields single-epoch and multi-epoch runs for the larger models but not for the smaller models (e.g. for the model with 85M parameters the maximum number of training tokens corresponds to 0.99 epochs). To better fit Equation 7, we performed additional experiments so that for each model size there were runs with training epochs in $\{2,4,8,10\}$, with the exception of the 828M parameter model, for which the maximum was 8 epochs.
@@ 4.1 Gains from sTinyStories
4.2 Benchmarking our setup
To validate our setup, we compared our best performing model with other models in the SLM literature in Table 3. Our model outperformed all other speech-only LMs on the semantic tasks, and performed second best in general, even relative to hybrid speech-text LMs. Notably, our model outperformed models with a larger compute budget. Considering that the models from Hassid et al. (2023) and Nguyen et al. (2024) use similar hyperparameters (same speech tokenizer and the Llama architecture for LMs); the most likely factor to explain the performance difference is the data used. We believe these results further illustrate the benefits from using sTinyStories.
Parameters Tokens BLIMP Topic Cloze Story Cloze
Speech-only language models
GSLM Lakhotia et al. (2021)  100M - 54.2 66.6 53.3
AudioLM Borsos et al. (2023)  150M - 64.7 - -
Hassid et al. (2023), Cold-init 1.3B  1.3B  10.8B 56.5 - -
Nguyen et al. (2024)  7B  100B 58.0 72.9 54.8
Ours (best model) 823M  82B 61.3 78.0 56.7
Speech language models initialized from text language models
TWIST Hassid et al. (2023)
   - Warm-init 1.3B  1.3B  10.8B 57.1 70.6 52.4
   - Warm-init 7B  7B  36B 59.0 74.1 55.1
   - Warm-init 13B  13B  36B 59.2 76.4 55.4
Mutltimodal speech-text language models initialized from text language models
SpiRit-LM Nguyen et al. (2024)  7B  100B 58.3 82.9 61.0
Toplines
Pythia Biderman et al. (2023) 6.9B  6.9B  300B 80.0 97.5 76.21
Human Hassid et al. (2023) - - - 90.2 79.9
Table 3: Models benchmarking. The best model resulting from our experiments obtains the best semantic performance across speech-only models, and the second best overall in all tasks.
@@ 4.3.1 Downstream scaling with compute
We analyzed the relationship between the upstream and linguistic downstream performance in SLMs and LLMs. Figure 4 shows the obtained results. Downstream linguistic metrics before saturation are strongly correlated with the upstream test loss in both LLMs and SLMs. Therefore, the envelope of maximum downstream performance per FLOP also follows a power law, i.e. for a downstream performance function $Q$, $Q\propto C^{\gamma_{q}}$. The power laws for the different performance metrics are presented in Figure 2 and the exponents in Table 4.
@@ 5 Related work
This paper is most closely related to the work of Hassid et al. (2023). We largely follow their setup in terms of hyperparameters and evaluation metrics. They reported improved linguistic downstream performance with scale in SLMs, but did not characterize their scaling behavior. Our scaling laws allow practitioners to determine the compute needed to attain a specific loss, syntactic and/or semantic performance; and its optimal allocation with respect to parameters and tokens. To the best of our knowledge, we are the first to model the scaling properties of downstream linguistic performance in SLMs, and to study the scaling of the considered downstream metrics on text-based LLMs. This enables a comparison between the two modalities in terms of scaling efficiency.
@@ 6 Discussion
7 Limitations
Any extrapolation from our models of the scaling behavior of SLMs should be considered optimistic for the following reasons: 1) Our models for downstream performance ignore the fact that the metrics saturate. As observed in text LLMs, the improvements with scale slow down as performance approaches the saturation value. It is likely that, due to saturation, the compute required to yield a particular performance will be larger than predicted. Moreover, due to the lower density of linguistic information per context window in SLMs relative to LLMs, the saturation values of the metrics may be lower for SLMs. 2) The LLMs from the Pythia suite that we used in this study are likely overtrained (all models were trained with $\sim$300B tokens). Optimally trained LLMs (according to Equation 6) should show better performance with scale, and therefore widen the gap with the scaling efficiency of SLMs. 3) The envelope of minimal loss per FLOP (Figure 1) might show a slight negative curvature at larger scale Hoffmann et al. (2022), reducing the scaling efficiency.
@@ 8 Conclusions
We have trained a large set of SLMs with different compute budgets and studied the scaling properties of their upstream and downstream performance using recently proposed models of scaling laws for neural LMs. The obtained models allow practitioners to optimally allocate compute to attain a specific loss, syntactic, and/or semantic performance. We showed that the pre-training loss and downstream linguistic performance of SLMs and LLMs is highly correlated, and both scale predictably according to power laws. This allowed us to compare the scaling properties of SLMs and LLMs, from which we established that the linguistic abilities of SLMs scale up to three orders of magnitude more slowly. Additionally, we proposed a new speech dataset, sTinyStories, and showed that its use during pre-training improves downstream semantic performance. Finally, we explored the use of coarser speech tokenization as a method to increase the amount of tokens per context window in SLMs, but obtained worse downstream performance.